OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 @satellite73 would this be no solution?

OpenStudy (anonymous):

no this one you can solve

OpenStudy (anonymous):

okay can you walk me through on how to solve this or similar? absolute values confuse me

OpenStudy (anonymous):

\[|2x+6|=16\] is a start then \[2x+6=16\] or \[2x+6=-16\] solve for \(x\)

OpenStudy (anonymous):

x=-5

OpenStudy (anonymous):

Quick hint: Always get your absolute values by themselves on one side of the equation. In this case, just add 4 to both sides to do that first.

OpenStudy (anonymous):

Sorry to interrupt @satellite73, I leave you to your excellent work.

OpenStudy (anonymous):

so 2x+6=24?

OpenStudy (anonymous):

oh no i made a huge mistake didn't i?

OpenStudy (anonymous):

\[|2x+6|-4=20\\ |2x+6|=24\] so \(2x+6=24\) or \(2x+6=-24\) forget my first post

OpenStudy (anonymous):

x=9 x=-15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

=))

OpenStudy (anonymous):

can you help me with 1 more?

OpenStudy (anonymous):

k